# Artificial Neural Networks for Magnetoencephalography: A review of an emerging field

### v1

@@ 3.3 Modeling studies
The idea of using AI techniques, particularly deep learning, to reverse engineer brain function has turned into a thriving topic in Neuro-AI. In the context of this review, we refer to ’Modeling’ studies that use ANNs to build models of some function of the brain. In contrast to the papers that use ANNs to classify MEG data, the studies in this category do not use MEG data as input for ANNs. Instead, they primarily involve comparisons between the ANN activations and MEG recordings obtained in response to presenting the same visual or auditory stimuli to both systems (i.e. the ANN and the human brain). Therefore, the difference between the MEG-ANN modeling work and the ANN-powered MEG classification work is fundamental and conceptual. The underlying rationale for conducting such comparisons is the hypothesis that higher similarities between the artificial and biological neural responses indicate greater functional similarities between the ANN model and the neural network mechanisms.
@@
Although the ANN training parameters, regularization, and validation techniques remain essential in this context for model training, they do not have the same type of impact on the results as for the classification studies discussed in the previous section. In the following, we’ll overview the main trends observed across the modeling studies, starting with the principle sensory processes that have been explored.
@@
Four out of the 16 studies in this category investigated the Auditory cortex ([98, 99, 100, 101]), while nine focus on the visual cortex ([87, 88, 89, 90, 91, 92, 93, 96, 97]). Additionally, two studies aimed to model the visual word recognition in the human brain ([94, 95]) and two set out to model speech recognition ([102, 103]). Out of the eleven studies investigating the visual cortex (including visual word recognition), eight use Representational Similarity Analysis (RSA) or a variant RSA-based approach, one uses a method based on ’neural predictivity’ ([92]), and the last two use ’profile responses’ ([94, 95]). Among the studies focusing on modeling the function of the auditory cortex (including speech recognition), two comparing representations of ANNs to those of the brain when presented with auditory stimuli also use a variant of RSA ([101, 103]). Three employ methods based on neural predictivity ([98, 99, 102]). The last study employs ’Dimensionality Analysis’ ([100]).
3.3.1 Multivariate Pattern Analysis (MVPA)
Nine out of the 16 studies surveyed in this category begin with a Multivariate Pattern Analysis (MVPA) analysis ([87, 88, 89, 90, 93, 94, 95, 96], [100]). MVPA is a statistical approach in neuroimaging that uses standard machine learning algorithms, such as LDA or SVM, to analyze and interpret patterns of brain activity across multiple voxels or sensors, thereby facilitating the identification of cognitive states from detailed and multifaceted neural data. But why do many studies in this category begin by using MVPA? The primary objective of conducting MVPA analysis before initiating the similarity analysis between MEG and ANN is to help identify patterns and features in the data most relevant for distinguishing between different mental states or stimuli. Applying MVPA first effectively reduces the dimensionality of the MEG data. It brings focus to the most informative features, which can then be used to compare stimulus representations across artificial and biological networks.
It is helpful to note here that out of the studies which used MVPA, three actually use MVPA results (i.e. decoding accuracy) to build similarity measures ([89, 90, 100]). Furthermore, some studies describe the entire process of classification followed by similarity analysis as MVPA.
3.3.2 Modeling pipeline
@@
Most of the articles included in the modeling category essentially compare neuromagnetic activity and the ANN activations generated by the same task (e.g., visual stimulus categorization).
@@
More concretely, once the models are trained, the same stimuli used for collecting brain data are fed to the ANNs and the layers’ responses’ are extracted. These responses are then considered as the ANNs’ activity, which will then be compared to the MEG activity induced by the same set of stimuli. That said, in one study, the authors concatenated the responses from all the layers responses to create a single model-level response [89]. Once the brain data are collected and the ANNs’ activations have been extracted, the next step is to compare them. In the following, we will review the main two methods used across the studies we surveyed: Representational Similarity Analysis (RSA) and neural predictivity. The other two remaining methods used in this context are ’Profile responses’ [94, 95] and ’Dimensionality analysis’ [100].
Representational Similarity Analysis (RSA): RSA is an established and widely used computational method to compare patterns of neural activity across different conditions or stimuli. This is achieved by constructing and analyzing similarity matrices representing the correlation or distance between the neural responses to each pair of conditions. These similarity matrices are known as Representational Dissimilarity Matrices (RDMs), which serve as a crucial tool in RSA by providing a quantitative measure of how neural responses differ across various experimental conditions, thus facilitating a deeper understanding of the underlying representational structures in the brain. In some studies, Representational Similarity Matrices (RSMs) are used instead of RDMs, employing direct correlation as the similarity metric rather than (1-correlation) for RDMs. Occasionally, decoding accuracies, an indirect measure of dissimilarity, are used instead of correlation measures ([89, 90, 100]).
@@
By contrast to RSA applications to fMRI data, the high temporal resolution of MEG provides the opportunity to conduct time-resolved RSA either in the sensor or in source space. To handle this additional time dimension, six of the 11 studies using RSA consider a 1D response vector for each time point by aggregating the activity of all brain sensors or specific ROIs’ voxels’ responses (in the case of source space). Ultimately, they generate time-resolved RDMs, sometimes referred to as ’RDM movies’ (see [91]). This particular RSA variant was termed RDA by [91]. Among the remaining studies employing RSA, two used a temporal variant from searchlight RSA, termed spatiotemporal sRSA (ssRSA) [101]. Instead of computing RDMs for every time point, the authors of these studies use time segments of a fixed length (25 ms for [101] and 16 ms for [87]). These studies combined this approach with searchlight RSA to construct multiple RDMs for different regions using a specific time interval. The final studies represented a combination of both approaches [93, 97, 103].
@@
Finally, a similarity score is computed by correlating the RDMs of MEG data with those of ANNs. The objective is to evaluate the similarity between the geometrical representations of MEG and ANNs across both time and space. For more information on RSA, see [139, 140, 141].
@@ 3.3.3 Estimating performance baselines
Assessing baseline performance is a crucial step in all analysis pipelines that seek to detect meaningful similarities between ANNs activity and MEG data. Generally speaking, determining baseline performance levels helps to distinguish the meaningfulness of the similarities observed between the ANN representations and the brain data, setting a benchmark for significance. Essentially, by establishing what can be expected by chance, or alternatively, what is maximally achievable given the noise in the data, we can confidently assert whether the similarities measured by RSA reflect genuine and meaningful correspondences between artificial and biological brain activity. In the following, we review the main methods of baseline performance used in the reviewed papers, focusing on Noise Ceilings (NC) and assessment of the untrained model’s performance.
@@
Noise Ceiling (NC), or the shared-response model (SRM) comparison, refers to the theoretical upper limit or maximum level of similarity that can be achieved in the absence of measurement noise or variability in brain data. In other words, a noise ceiling provides an estimate of the best performance any model can achieve given the noise in the data. As such, it serves as a benchmark for the observed similarities between ANNs and the brain, indicating the extent to which individual subjects’ brain responses can be explained with a model-free approach. This concept often acts as a proxy for signal-to-noise ratio analysis, aiding researchers in interpreting the significance and reliability of their observed similarities by considering the inherent noise or variability in the data. As is common in RSA, the upper noise ceiling is estimated as the mean correlation between the group-average RSM and each participant-specific RSM. The lower noise ceiling is estimated as the mean correlation between the group-average RSM and each participant-specific RSM while iteratively excluding a given participant from the group-average. Some studies report both upper and lower noise ceilings. Although it is an essential measure, only five of the surveyed studies reported it. Among them, four used RSA as a similarity method ([87, 88, 91, 97]), one used neural predictivity ([99]) and the last used profile responses ([94, 95]). For detailed instructions on how to compute the noise ceiling, refer to [88].
@@
Untrained models performance as baseline In RSA studies, a common approach to assess performance is to compare the results obtained with the trained ANN model to those obtained with an untrained ANN model. Some RSA-based studies (e.g. [143]) have found that even randomly initialized models can exhibit some similarity to the neural representation. Such an intrinsic similarity could result from built-in properties of specific models’ architecture (e.g., convolutional layers resembling the visual cortex). However, most RSA studies make implicit assumptions (or at least have some expectations) regarding the fact that trained ANNs would lead to enhanced ANN-brain similarities. In this context, comparisons of RSA results obtained with trained and untrained models can be very informative. Interestingly, out of the sixteen studies we found in this category, only three contrast their modeling results with those of a random model ([89, 94, 95, 99]).
@@
While most studies focus on a single architecture, some include multiple ANN models. By conducting RSA analyses between the brain data and each one of the models, such studies can pick up the specific network architecture and training properties that increase the similarities between the ANN and biological responses. In [87], the authors employed a feature-based model where RDMs are computed based on the features extracted from the stimuli. In [89], the authors added a model trained on noise and an unecological model where images have been assigned random labels. Furthermore, in [88], the authors compared their similarities with those obtained by a semantic model. In [97], the authors use semantic models that measure semantic similarity between objects, i.e., the degree of resemblance in meaning between two pieces of text, such as words or sentences. In addition, it is common to train the same ANN on several similar tasks to assess the importance of the objective functions.
3.3.4 The data
Before we discuss the data used in the modeling papers, it is essential to recall that, unlike studies that use ANNs to classify MEG data, the research reviewed in this category does not input MEG data into ANNs. Instead, it primarily compares ANN activations with MEG recordings, observing responses to the same visual or auditory stimuli presented separately to both systems. For instance, in the study by Kietzmann et al. [91], the input data consists of images from different categories (animate or inanimate objects, faces, etc). Although MEG data are not used as inputs to the ANNs in these studies, they are still used to compare the representations in brain signals and the latent space variables of ANNs.
All MEG data collected for the studies in this category were preprocessed using a similar procedure, which in principle consists of a band-pass filter, an artifact removal and/or correction technique, and data down-sampling. The band-pass filtering’s lower and upper cutoff frequencies are 0.03 or 0.1Hz to 300 or 330Hz. However, there seems to be no discernible pattern explaining the sampling frequencies, the de-noising, or artifact removal techniques used (see 5). Among the 10 studies using RSA, five have used source reconstruction before computing RDMs ([87, 91, 101, 94, 95, 97]). It is worth noting that only two studies explored the frequency domain of MEG data by investigating the similarities in one or more frequency bands ([100, 99]).
3.3.5 The networks
The choice of ANN architecture is a particularly relevant aspect of the ANN-MEG modeling studies because the ultimate goal here is to assess and interpret similarities and discrepancies between the way information is processed in biological and artificial networks. As a general rule, the articles that focus on the visual cortex use a CNN architecture trained on image classification. Out of these studies, one incorporates lateral connections between layers, in addition to the typical top-down connections, introducing temporal relationships in their ANN ([91]). In this study, the ANN is trained to learn how to recreate MEG data RDMs from the stimuli image displayed to the subject. This is achieved through changing the ANN’s objective function. About half of the studies interested in the auditory cortex use LSTM ([100, 98]), the remaining studies use Transformer network and a CNN ([99]), an MLP architecture ([101]) or the BERT ([144]) model [103]. Finally, in [94, 95], the authors investigated the neural responses for a visual word recognition task by comparing the visual cortex’s response patterns to those of CNN architectures.

### v3

@@ 3.3.1 Study aims and subcategories
The idea of using AI techniques, particularly deep learning, to reverse engineer brain function has turned into a thriving topic in Neuro-AI. In the context of this review, we refer to ’Modeling’ studies that use ANNs to build models of some function of the brain. In contrast to the papers that use ANNs to classify MEG data, the studies in this category do not use MEG data as input for ANNs. Instead, they primarily involve comparisons between the ANN activations and MEG recordings obtained in response to presenting the same visual or auditory stimuli to both systems (i.e. the ANN and the human brain). Therefore, the difference between the MEG-ANN modeling work and the ANN-powered MEG classification work is basic. The underlying rationale for conducting such comparisons is the hypothesis that higher similarities between the artificial and biological neural responses indicate greater functional similarities between the ANN model and the neural network mechanisms.
@@
Although the ANN training parameters, regularization, and validation techniques remain essential in this context for model training, they do not have the same type of impact on the results as for the classification studies discussed in the previous section. In the following, we will overview the main trends observed across the modeling studies, starting with the principle sensory processes that have been explored.
Four out of the 16 studies in this category investigated the auditory cortex ([128, 129, 131, 130]), while nine focused on the visual cortex ([121, 122, 118, 119, 124, 120, 125, 123, 127]). Additionally, one study aimed to model the visual word recognition in the human brain ([126]) and two set out to model speech recognition ([132, 133]).
Out of the eleven studies investigating the visual cortex (including visual word recognition), eight use representational similarity analysis (RSA [185]) or a variant RSA-based approach, one used a method based on ’neural predictivity’ ([120]), and the last one used ’profile responses’ ([126]).
@@
Among the studies focusing on modeling the function of the auditory cortex (including speech recognition), two comparing representations of ANNs to those of the brain when presented with auditory stimuli also used a variant of RSA ([130, 133]). Three employed methods based on neural predictivity ([128, 129, 132]). The last study employed ’dimensionality analysis’ ([131]).
3.3.2 The pipeline
@@
A schematic overview of a typical pipeline for studies in the ’Modeling’ category is provided in figure 4 (center panel). Most of the articles included in this category essentially compare neuromagnetic activity and the ANN activations generated by the same task (e.g., visual stimulus categorization).
@@
More concretely, once the models are trained, the same stimuli used for collecting brain data are fed to the ANNs and the layers’ responses’ are extracted. These responses are then considered as the ANNs’ activity, which will then be compared to the MEG activity induced by the same set of stimuli. That said, in one study, the authors concatenated the responses from all the layers responses to create a single model-level response ([118]). Once the brain data are collected and the ANNs’ activations have been extracted, the next step is to compare them. In the following, we will review the main two methods used across the studies we surveyed: RSA and neural predictivity. The remaining method used in this context are ’profile responses’ ([126]) and ’dimensionality analysis’ ([131]).
Representational similarity snalysis: RSA is an established and widely used computational method to compare patterns of neural activity across different conditions or stimuli. This is achieved by constructing and analyzing similarity matrices representing the correlation or distance between the neural responses to each pair of conditions. These similarity matrices are known as representational dissimilarity matrices (RDMs), which serve as a crucial tool in RSA by providing a quantitative measure of how neural responses differ across various experimental conditions, thus facilitating a deeper understanding of the underlying representational structures in the brain. In some studies, representational similarity matrices (RSMs) are used instead of RDMs, employing direct correlation as the similarity metric rather than $(1-\mathrm{correlation})$ for RDMs. Occasionally, decoding accuracies, an indirect measure of dissimilarity, are used instead of correlation measures ([118, 119, 131]).
@@
By contrast to RSA applications to fMRI data, the high temporal resolution of MEG provides the opportunity to conduct time-resolved RSA either in the sensor or in source space. To handle this additional time dimension, six of the 11 studies using RSA consider a 1D response vector for each time point by aggregating the activity of all brain sensors or specific ROIs’ voxels’ responses (in the case of source space). Ultimately, they generate time-resolved RDMs, sometimes referred to as ’RDM movies’ (see [124]). This particular RSA variant was termed RDA by [124]. Among the remaining studies employing RSA, two used a temporal variant from searchlight RSA, termed spatiotemporal sRSA (ssRSA [130]). Instead of computing RDMs for every time point, the authors of these studies use time segments of a fixed length (25 ms for [130] and 16 ms for [121]). These studies combined this approach with searchlight RSA to construct multiple RDMs for different regions using a specific time interval. The final studies represented a combination of both approaches ([125, 127, 133]).
@@
Finally, a similarity score is computed by correlating the RDMs of MEG data with those of ANNs. The objective is to evaluate the similarity between the geometrical representations , defined as neural activity patterns organized into a geometric space according to the distances among these patterns, of MEG and ANNs. This comparison is performed across both time (how MEG neural patterns evolve and align temporally with ANN activations) and space (how spatial patterns from MEG sensors or sources relate to ANN activations). Such comparison reveals whether and when ANNs produce neural representations resembling those of the human brain. For more information on RSA, see [186, 187, 185].
@@
Multivariate pattern analysis (MVPA): Nine out of the 16 studies surveyed in this category begin with an MVPA analysis ([121, 122, 118, 119, 125, 126, 123, 131]). MVPA is a statistical approach in neuroimaging that uses standard machine learning algorithms, such as LDA and SVM, to analyze and interpret patterns of brain activity across multiple voxels or sensors, thereby facilitating the identification of cognitive states from thorough and varied neural data. But why do many studies in this category begin by using MVPA? The primary objective of conducting MVPA analysis before initiating the similarity analysis between MEG and ANN is to help identify patterns and features in the data most relevant for distinguishing between different mental states or stimuli. Applying MVPA first effectively reduces the dimensionality of the MEG data. It brings focus to the most informative features, which can then be used to compare stimulus representations across artificial and biological networks.
It is helpful to note here that out of the studies which used MVPA, three actually used MVPA results (i.e. decoding accuracy) to build similarity measures ([118, 119, 131]). Furthermore, some studies described the entire process of classification followed by similarity analysis as MVPA.
3.3.3 The data and preprocessing
Across the 16 modeling studies reviewed, the number of participants ranged from 11 to 92 (median = 15), and the number of trials or samples ranged from 77 to 66,300 (median = 3,585). However, five studies did not report the number of samples, and none reported the number of trials used for model training explicitly, as these studies did not train ANNs on MEG data.
It is important to note that, unlike classification studies where MEG data serve as input to ANNs, modeling studies in this review did not train ANNs using MEG signals. Instead, it primarily compares ANN activations with MEG recordings, observing responses to the same visual or auditory stimuli presented separately to both systems. For instance, in the study by Kietzmann et al. ([124]), the input data consists of images from different categories (animate or inanimate objects, faces, etc). Although MEG data are not used as inputs to the ANNs in these studies, they are still used to compare the representations in brain signals and the latent space variables of ANNs.
All MEG data collected for the studies in this category were preprocessed using a similar procedure, which in principle consists of a band-pass filter, an artifact removal and/or correction technique, and data down-sampling. The band-pass filtering’s lower and upper cutoff frequencies are 0.03 or 0.1 Hz to 300 or 330 Hz. However, there seems to be no discernible pattern explaining the sampling frequencies (figure 3), the de-noising, or artifact removal techniques used (see table 4). Among the 10 studies using RSA, five have used source reconstruction before computing RDMs ([121, 124, 130, 126, 127]). It is worth noting that only two studies explored the frequency domain of MEG data by investigating the similarities in one or more frequency bands ([131, 129]).
3.3.4 Network architectures
The choice of ANN architecture is a particularly relevant aspect of the ANN-MEG modeling studies because the ultimate goal here is to assess and interpret similarities and discrepancies between the way information is processed in biological and artificial networks. As a general rule, the articles that focus on the visual cortex use a CNN architecture trained on image classification. Out of these studies, one incorporates lateral connections between layers, in addition to the typical top-down connections, introducing temporal relationships in their ANN ([124]). In this study, the ANN is trained to learn how to recreate MEG data RDMs from the stimuli image displayed to the subject. This is achieved through changing the ANN’s objective function. About half of the studies interested in the auditory cortex use LSTM ([131, 128]), the remaining studies use transformer network and a CNN ([129]), an MLP architecture ([130]) or the BERT ([189]) model ([133]). Finally, in [126], the authors investigated the neural responses for a visual word recognition task by comparing the visual cortex’s response patterns to those of CNN architectures.
3.3.5 Training and validation techniques
In modeling studies, the emphasis is placed less on predictive performance and more on the alignment between ANN representations and brain responses. As a result, training and validation procedures are not always detailed with the same rigor as in classification-focused work. Still, when custom architectures are trained, typical practices include specifying the optimization algorithm, loss function, and stopping criteria, though many studies using pre-trained networks omit these details entirely.
Among the studies that reported their training procedures, most adopted standard deep learning practices such as stochastic optimization and loss minimization over a supervised objective. The Adam optimizer was frequently employed due to its computational efficiency and robustness to noisy gradients, although few papers provided full training specifications. In modeling studies, ANNs are typically pretrained or fine-tuned on tasks unrelated to MEG data though these tasks are often similar to the experimental conditions used to collect the MEG data (e.g., object recognition, language processing). Training and validation procedures focus on the model’s primary objective (e.g., classification or language modeling), while representational alignment with MEG signals is assessed in a separate analysis using similarity metrics such as RSA. These approaches differ fundamentally from classification pipelines, as MEG data is not used to train or validate the model itself, but rather to evaluate how well its internal representations reflect brain activity.
@@ 3.3.6 Estimating performance baselines
Assessing baseline performance is crucial for benchmarking and evaluating the significance of the observed similarities. Essentially, by establishing what can be expected by chance, or alternatively, what is maximally achievable given the noise in the data, we can confidently assert whether the similarities measured by RSA reflect genuine and meaningful correspondences between the representations in artificial neural networks and those observed in biological neural activity. In the following, we review the main methods of baseline performance used in the reviewed papers, focusing on noise ceilings (NC) and assessment of the untrained model’s performance.
@@
Noise ceiling, or the shared-response model (SRM) comparison, refers to the theoretical upper limit or maximum level of similarity that can be achieved in the absence of measurement noise or variability in brain data. In other words, a noise ceiling provides an estimate of the best performance any model can achieve given the noise in the data. As such, it serves as a benchmark for the observed similarities between ANNs and the brain, indicating the extent to which individual subjects’ brain responses can be explained with a model-free approach. This concept often acts as a proxy for signal-to-noise ratio analysis, aiding researchers in interpreting the significance and reliability of their observed similarities by considering the inherent noise or variability in the data. As is common in RSA, the upper noise ceiling is estimated as the mean correlation between the group-average RSM and each participant-specific RSM. The lower noise ceiling is estimated as the mean correlation between the group-average RSM and each participant-specific RSM while iteratively excluding a given participant from the group-average. Some studies report both upper and lower noise ceilings. Although it is an essential measure, only five of the surveyed studies reported it. Among them, four used RSA as a similarity method ([121, 122, 124, 127]), one used neural predictivity ([129]) and the last used profile responses ([126]). For detailed instructions on how to compute the noise ceiling, refer to [122].
@@
Untrained models performance as baseline: In RSA studies, a common approach to assess performance is to compare the results obtained with the trained ANN model to those obtained with an untrained ANN model. Some RSA-based studies (e.g. [191]) have found that even randomly initialized models can exhibit some similarity to the neural representation. Such an intrinsic similarity could result from built-in properties of specific models’ architecture (e.g., convolutional layers resembling the visual cortex). However, most RSA studies make implicit assumptions (or at least have some expectations) that training ANNs would lead to enhanced ANN-brain similarities. In this context, comparisons of RSA results obtained with trained and untrained models can be very informative. Interestingly, out of the sixteen studies we found in this category, only three contrast their modeling results with those of a random model ([118, 126, 129]).
@@
While most studies focus on a single architecture, some include multiple ANN models. By conducting RSA analyses between the brain data and each one of the models, such studies can pick up the specific network architecture and training properties that increase the similarities between the ANN and biological responses. In [121], the authors employed a feature-based model where RDMs are computed based on the features extracted from the stimuli. In [118], the authors added a model trained on noise and an unecological model where images have been assigned random labels. Furthermore, in [122], the authors compared their similarities with those obtained by a semantic model. In [127], the authors use semantic models that measure semantic similarity between objects, i.e., the degree of resemblance in meaning between two pieces of text, such as words or sentences. In addition, a new trend consists of training the same ANN on several related tasks (or use different objective functions) to investigate how each training goal shapes the learned representations ([192, 193, 194]).